### v1

# Periodic Boundary Conditions for Bosonic Path Integral Molecular Dynamics

Abstract
We develop an algorithm for bosonic path integral molecular dynamics (PIMD) simulations with periodic boundary conditions (PBC) that scales quadratically with the number of particles. Path integral methods are a powerful tool to simulate bosonic condensed phases, which exhibit fundamental physical phenomena such as Bose–Einstein condensation and superfluidity. Recently, we developed a quadratic scaling algorithm for bosonic PIMD, but employed an ad hoc treatment of PBC. Here we rigorously enforce PBC in bosonic PIMD. It requires summing over the spring energies of all periodic images in the partition function, and a naive implementation scales exponentially with the system size. We present an algorithm for bosonic PIMD simulations of periodic systems that scales only quadratically. We benchmark our implementation on the free Bose gas and a model system of cold atoms in optical lattices. We also study an approximate treatment of PBC based on the minimum-image convention, and derive a numerical criterion to determine when it is valid.
I Introduction
Path integral molecular dynamics (PIMD) is a prominent method for calculating properties of quantum condensed phases at thermal equilibrium Parrinello and Rahman (1984); Tuckerman et al. (1993). In PIMD, thermal expectation values of the quantum system are inferred from molecular dynamics simulations of an extended classical system of “ring polymers”. The ring polymers are formed by connecting $P$ copies of a quantum particle (“beads” or “imaginary time slices”) through harmonic springs whose stiffness depends on the temperature. PIMD simulations are widely used to study quantum liquids and solids Markland and Ceriotti (2018). They are also the basis for several methods for approximating quantum transport coefficients Althorpe (2021).
Recently, we developed PIMD simulations of bosonic systems. Hirshberg et al. (2019); Feldman and Hirshberg (2023) The challenge was that bosonic exchange symmetry required summing over an exponential number of ring polymer configurations in the partition function, each formed by connecting the rings of exchanged particles together. We overcame this combinatorial explosion, and presented an algorithm that scales quadratically with the size of the system, $\mathcal{O}(N^{2}+PN)$, where $N$ is the number of particles and $P$ is the number of beads per particle. This quadratic scaling allows efficient simulations of thousands of bosons Feldman and Hirshberg (2023). This method is the basis also for several techniques studying fermionic systems Hirshberg et al. (2020); Dornheim et al. (2020); Xiong and Xiong (2022a, b, c); Xiong et al. (2024).
An important tool in studying quantum condensed phases is the use of periodic boundary conditions (PBC) to capture properties of the the bulk through the simulation of a smaller system Allen and Tildesley (2017). However, previous applications of bosonic PIMD have focused primarily on trapped, non-periodic systems, and the occasional periodic system was handled by ad hoc methods (as we explain in Section IV). The goal of this paper is to develop an efficient, rigorous bosonic PIMD algorithm for periodic systems.
How should the PBC modify bosonic path integral simulations? To properly account for PBC, the partition function should account for the spring interaction of every bead with all periodic images of its neighboring beads Ceperley (1995); Marx and Müser (1999); Kleinert (2009). The challenge is that there are exponentially many different winding configurations that need to be taken into account in the simulation. In path integral Monte Carlo (PIMC), the problem is addressed by sampling configurations in which the springs are stretched or compressed by integer multiples of the box length, referred to as the winding of the springs Cao (1994); Marx et al. (1993); Miller and Clary (2002); Spada et al. (2022); Del Maestro et al. (2022). In practice, the sampling of the windings is done up to some cutoff $\mathcal{W}$, because configurations with highly stretched springs are energetically improbable.
The primary contribution of this paper is the development of an efficient bosonic PIMD algorithm with PBC that scales as $\mathcal{O}(\mathcal{W}(N^{2}+PN))$. While in PIMC a winding configuration for the path is sampled at each step, in PIMD it is necessary to consider all possible winding configurations simultaneously. Furthermore, winding and bosonic exchange are coupled, i.e., the partition function cannot be decomposed into independent sums over windings and permutations since the spring energy depends non-linearly on both. Despite these challenges, we present an algorithm for periodic bosonic PIMD that has the same quadratic scaling with system size as the previous algorithm, which did not include summation over windings. The new algorithm also has linear scaling with the winding cutoff $\mathcal{W}$. Thus, our algorithm rigorously extends bosonic PIMD to periodic systems.
In Section II we present the theoretical background for the paper. Sections II.1 and II.2 summarize the theory of bosonic PIMD with quadratic scaling without accounting for the windings. Section II.3 presents the partition function with PBC, and explains the challenges in sampling it. Section III presents the key results of the paper. First, Section III.1 explains how to perform PIMD simulations with PBC for the case of distinguishable particles. Then, we explain the theory behind the new bosonic algorithm in Section III.2, and in Section III.3, we benchmark it on two model systems: free Bose gas and a model of cold bosons in an optical lattice. In Section IV, we rigorously examine a simpler approximation for PBC: applying the minimum-image convention (MIC) to the springs. We develop a quantitative criterion to when MIC can be used instead of the windings algorithm.
II Background
II.1 Bosonic PIMD
In non-periodic systems, the path integral expression for the canonical partition function of $N$ bosons at thermal equilibrium with inverse temperature $\beta$, in the presence of the physical potential $U$, is Tuckerman (2010)
[…]
In Equation 1, ${\bf R}_{\ell}={\bf r}_{\ell}^{1},\ldots,{\bf r}_{\ell}^{P}$ collectively represents the position vectors of all the $j=1,...,P$ beads of particle $\ell$, and the expression is exact in the limit $P\to\infty$. The sum in Equation 1 is over all permutations of the $N$ bosons. Each permutation $\sigma$ corresponds to a ring polymer configuration in which particles are connected according to the permutation, i.e., the last bead of particle $\ell$ is connected to the first bead of particle $\sigma(\ell)$. The spring energy of a configuration is
$E^{\sigma}=\frac{1}{2}m\omega_{P}^{2}\sum_{\ell=1}^{N}{\sum_{j=1}^{P}{\left({\bf r}_{\ell}^{j}-{\bf r}_{\ell}^{j+1}\right)^{2}}},$ (2)
with ${\bf r}_{\ell}^{P+1}={\bf r}_{\sigma(\ell)}^{1}$, and $\omega_{P}={\sqrt{P}}/{(\beta\hbar)}$. Beads $j$ of different particles interact according to the scaled potential $\bar{U}=\frac{1}{P}\sum_{j=1}^{P}{U\left({\bf r}_{1}^{j},\ldots,{\bf r}_{N}^{j}\right)}$, which is invariant under particle permutations.
II.2 Quadratic scaling algorithm for bosonic PIMD
Due to the sum over an exponential number of permutations, directly sampling the partition function through Equation 1 is computationally prohibitive. In previous work Hirshberg et al. (2019); Feldman and Hirshberg (2023), we showed that the same partition function can be expressed in a way that is amenable to efficient computation, by writing
[…]
where the bosonic spring potential $V^{[1,{N}]}$ is defined by the recurrence relation
$e^{-\beta{V^{[1,{N}]}}}$ $=\frac{1}{N}\sum_{k=1}^{N}{e^{-\beta{\left(V^{[1,{N-k}]}+E^{[{{N}-{k}+1},{{N}}]}\right)}}}.$ (4)
The recurrence is terminated by $V^{[1,{0}]}=0$. In Equation 4, $E^{[{{N}-{k}+1},{{N}}]}$ is the spring energy of the ring polymer that connects all the beads of particles $N-k+1,\ldots,N$ consecutively,
$E^{[{{N}-{k}+1},{{N}}]}=\frac{1}{2}m\omega_{P}^{2}\sum_{\ell=N-k+1}^{N}{\sum_{j=1}^{P}{\left({\bf r}_{\ell}^{j+1}-{\bf r}_{\ell}^{j}\right)^{2}}},$ (5)
where ${\bf r}_{\ell}^{P+1}={\bf r}_{\ell+1}^{1}$ except ${\bf r}_{N}^{P+1}={\bf r}_{N-k+1}^{1}$.
The potential $V^{[1,{N}]}$ allows to perform PIMD simulations to sample the bosonic partition function in polynomial scaling Hirshberg et al. (2019). We showed Feldman and Hirshberg (2023) that both the potential and the forces can be computed in quadratic time, $\mathcal{O}(N^{2}+PN)$. First, the bosonic spring potential is computed, by evaluating the quantities $E^{[{{N}-{k}+1},{{N}}]}$ through another recurrence relation, extending cycles one particle at a time. Then, the spring forces on all the beads are computed using expressions for the probabilities of the different ways particles can be connected. The full details of the previous algorithm appear in Ref. 6.
These equations do not rigorously address PBC. Our central achievement in this paper is developing an efficient PIMD method with quadratic scaling for periodic bosonic systems. Below, we first review the partition function with PBC, and, in Section III, derive an efficient PIMD algorithm to sample it.
II.3 Periodic boundary conditions
Pollock and Ceperley Pollock and Ceperley (1987) showed that imposing PBC alters the path integral expression for the partition function, replacing Equation 1 by
[…]
where $D(\mathcal{V})$ is the spatial domain defined by the volume of the unit cell. Throughout this paper, we assume a cubic box with side length $L$. In Equation 6, the energy of a configuration is
$E^{\sigma,\{{\mathbf{w}}\}}=\frac{1}{2}m\omega_{P}^{2}\sum_{\ell=1}^{N}{\sum_{j=1}^{P}{\left({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1}\right)^{2}}},$ (7)
with ${\bf r}_{\ell}^{P+1}={\bf r}_{\sigma(\ell)}^{1}$.
Figure 1: Winding configurations for $N=2$ particles in one-dimension with $P=2$ beads (blue for ${\bf r}_{1}^{j}$ and red for ${\bf r}_{2}^{j}$) each and winding cutoff of $\mathcal{W}=2$. In this case, there are only two permutations. The configurations differ by the winding of the spring connecting ${\bf r}_{1}^{P}$ to ${\bf r}_{\sigma(1)}^{1}$ (red solid line), where $\sigma$ is the permutation.
Equation 6 differs from Equation 1 in the additional sum over winding vectors. We denote the $d$-dimensional winding vector of bead $j$ of particle $\ell$ by $\mathbf{w}_{\ell}^{j}$. Its components are integers expressing how many times the spring that connects a bead to its next neighbor winds around the box along a certain axis (see the top row of Figure 1 for an example with $d=1$). The set of windings $\{{\mathbf{w}}\}=\mathbf{w}_{1}^{1},\ldots,\mathbf{w}_{1}^{P},\ldots,\mathbf{w}_{N}^{1},\ldots,\mathbf{w}_{N}^{P}$ represents a specific combination of integer components of the winding vectors of all the beads. Additionally, particles are connected to rings according to the permutation $\sigma$ as in Section II.1. Figure 1 depicts winding configurations of one of the springs for the two permutations of two bosons.
With this notation, the sum over $\{{\mathbf{w}}\}$ in Equation 6 is to be understood as
$\sum_{\{{\mathbf{w}}\}}=\sum_{\mathbf{w}_{1}^{1}}\dots\sum_{\mathbf{w}_{1}^{P}}\dots\sum_{\mathbf{w}_{N}^{1}}\dots\sum_{\mathbf{w}_{N}^{P}},$ (8)
where,
$\sum_{\mathbf{w}_{\ell}^{j}}\equiv\sum_{\left({w}_{\ell}^{j}\right)_{1}=-\mathcal{W}}^{\mathcal{W}}\dots\sum_{\left({w}_{\ell}^{j}\right)_{d}=-\mathcal{W}}^{\mathcal{W}}.$ (9)
In practice, the sum over all possible integer components of $\mathbf{w}_{\ell}^{j}$ in Equation 9 is capped Del Maestro et al. (2022) by introducing a non-negative integer cutoff $\mathcal{W}$. This is done because large windings are unfavorable as they result in large spring energies, per Equation 7.
The challenge for PIMD is that the number of configurations in the sum of Equation 6 is exponential in both the number of particles and the number of beads, $N!\cdot(2\mathcal{W}+1)^{dPN}$. The $N!$ term originates from particle exchange, and $(2\mathcal{W}+1)^{dPN}$ from the sum over windings. These two phenomena are intertwined because the spring energy of the last bead $j=P$ of each particle $\ell$ depends non-linearly on both the winding vector $\mathbf{w}_{\ell}^{P}$ and the permutation $\sigma(\ell)$ (see Figure 1). For this reason, it is impossible to apply the previous algorithm to mitigate the combinatorial explosion from particle exchange, and then handle PBC separately.
In this paper, we show how to efficiently include PBC in bosonic PIMD. Our algorithm scales quadratically with system size and only linearly with the winding cutoff, $\mathcal{O}(\mathcal{W}(N^{2}+PN))$. Compared to the previous algorithm, the new algorithm requires changes in all its stages: the definition of the recursive bosonic potential, its evaluation, and the force calculation. Before we explain the algorithm in Section III.2, we first present in Section III.1 an algorithm that efficiently includes PBC for the case of distinguishable particles.
III Results
III.1 Linear scaling of distinguishable PIMD with PBC
We first describe an efficient, $\mathcal{O}(\mathcal{W}PN)$, algorithm for distinguishable particles with PBC. The bosonic algorithm in Section III.2 builds on the ideas presented here and extends them to include particle exchange.
The partition function for distinguishable particles with PBC is obtained by including only the identity permutation in Equation 6,
[…]
through a distinguishable spring potential that includes only summation over windings,
$e^{-\beta{V_{\text{D}}}}=\sum_{\{{\mathbf{w}}\}}{e^{-\beta{E^{\{{\mathbf{w}}\}}}}}.$ (11)
The spring energy of a configuration is then
$E^{\{{\mathbf{w}}\}}=\frac{1}{2}m\omega_{P}^{2}\sum_{\ell=1}^{N}{\sum_{j=1}^{P}{\left({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1}\right)^{2}}},$ (12)
and ${\bf r}_{\ell}^{P+1}={\bf r}_{\ell}^{1}$. Note that PBC introduce, in principle, a sum over exponentially many winding configurations while, in the case of non-periodic distinguishable systems, there is only a single configuration. We will now show how to perform PIMD simulations of distinguishable particles with PBC in linear time in two steps: evaluating the potential, and evaluating the forces on all the beads.
III.1.1 Computing the potential in $\mathcal{O}(\mathcal{W}PN)$ time
We denote the statistical weight of a single spring with a specific winding by
$\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1},\mathbf{w}_{\ell}^{j})=e^{-\beta{\frac{1}{2}m\omega_{P}^{2}\left({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1}\right)^{2}}}.$ (13)
Then, the potential defined by Equations 11 and 12 can be written as a sum over products of individual weights.
$e^{-\beta{V_{\text{D}}}}=\sum_{\{{\mathbf{w}}\}}{\prod_{\ell=1}^{N}{\prod_{j=1}^{P}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1},\mathbf{w}_{\ell}^{j})}}}.$ (14)
Figure 2: An illustration of calculating the distinguishable potential using Equation 15 in the case of a single particle in one-dimension with $P=2$ beads and a winding cutoff of $\mathcal{W}=1$. In this case, there are $\left(2\mathcal{W}+1\right)^{dNP}=9$ different winding configurations surrounded by the dashed line. These configurations are constructed by first summing over windings of individual springs, represented by the “partial” configurations (the contribution from shaded areas are excluded) at the top row and left column, and then multiplying.
Because $\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1},\mathbf{w}_{\ell}^{j})$ depends only on a single winding, we can rewrite the sum of products over the set $\{{\mathbf{w}}\}$ as a product of sums over an individual winding vector $\mathbf{w}_{\ell}^{j}$
$e^{-\beta{V_{\text{D}}}}=\prod_{\ell=1}^{N}{\prod_{j=1}^{P}{\left(\sum_{\mathbf{w}_{\ell}^{j}}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1},\mathbf{w}_{\ell}^{j})}\right)}}.$ (15)
Figure 2 illustrates how the total contribution of all winding vectors to $V_{D}$ can be computed by first summing over individual winding vectors and then multiplying the resulting contributions across different beads. This expression shows that we evaluate the potential in $\mathcal{O}(\mathcal{W}PN)$ time: the contribution of each of the $PN$ springs involves a sum over its windings, of which there are $12\mathcal{W}+$ 1 possibilities. In more than one dimension, the sum over winding vectors of each spring in Equation 15 splits to the product of sums over each coordinate separately, hence the scaling is $\mathcal{O}(\mathcal{W})$ and not $\mathcal{O}(\mathcal{W}^{d})$—see the supplementary information (SI) in Section I.D for details. In the case of $\mathcal{W}=0$, the scaling of the algorithm reduces to the standard $\mathcal{O}(PN)$ distinguishable PIMD.
To simplify notation in the force derivation in the rest of the paper, we denote the total weight of all the windings of a given spring by
$\mu({\bf r}_{u}^{j},{\bf r}_{v}^{k})=\sum_{\mathbf{w}_{u}^{j}}{\mu({\bf r}_{u}^{j},{\bf r}_{v}^{k},\mathbf{w}_{u}^{j})},$ (16)
which leads to the following expression for the distinguishable spring potential
$e^{-\beta{V_{\text{D}}}}=\prod_{\ell=1}^{N}{\prod_{j=1}^{P}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{j+1})}}.$ (17)
III.1.2 Computing the forces in $\mathcal{O}(\mathcal{W}PN)$ time
We now turn to an efficient evaluation of the forces. To keep the reasoning as close as possible to the bosonic case, we start by taking the derivative of Equation 11 to see that the force can be written as a weighted average over windings:
$-\nabla_{{\bf r}_{\ell}^{j}}{V_{\text{D}}}={\sum_{\{{\mathbf{w}}\}}{\Pr({\{{\mathbf{w}}\}})\cdot-\nabla_{{\bf r}_{\ell}^{j}}{E^{\{{\mathbf{w}}\}}}}}.$ (18)
In Equation 18, the force exerted on a bead in a configuration with windings $\{{\mathbf{w}}\}$ depends only on two winding vectors:
$-\nabla_{{\bf r}_{\ell}^{j}}{E^{\{{\mathbf{w}}\}}}=$ $-m\omega_{P}^{2}({\bf r}_{\ell}^{j}-{\bf r}_{\ell}^{j-1}-\mathbf{w}_{\ell}^{j-1}L)$ (19)
$-m\omega_{P}^{2}({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1}).$
Additionally, $\Pr({\{{\mathbf{w}}\}})$ is a Boltzmann probability distribution over the configurations,
$\Pr({\{{\mathbf{w}}\}})=\frac{e^{-\beta{E^{\{{\mathbf{w}}\}}}}}{e^{-\beta{V_{\text{D}}}}}.$ (20)
To efficiently evaluate the force in Equation 18, we combine the contributions that exert the same force,
$-\nabla_{{\bf r}_{\ell}^{j}}{V_{\text{D}}}=$ $\sum_{\mathbf{w}_{\ell}^{j-1}}{\Pr\left(\mathbf{w}_{\ell}^{j-1}\right)\cdot-m\omega_{P}^{2}({\bf r}_{\ell}^{j}-{\bf r}_{\ell}^{j-1}-\mathbf{w}_{\ell}^{j-1}L)}$ (21)
$+$ $\sum_{\mathbf{w}_{\ell}^{j}}{\Pr\left(\mathbf{w}_{\ell}^{j}\right)\cdot-m\omega_{P}^{2}({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1})}.$
In Equation 21, the probability $\Pr\left(\mathbf{w}_{\ell}^{j}\right)$ is the sum of the probabilities of all configurations in which the winding of bead $j$ of particle $\ell$ is $\mathbf{w}_{\ell}^{j}$. If $\Pr\left(\mathbf{w}_{\ell}^{j-1}\right),\Pr\left(\mathbf{w}_{\ell}^{j}\right)$ are known for all $\ell$ and $j$, the force can be evaluated simply by summing over a single winding vector for each of the two spring force terms, in $\mathcal{O}(\mathcal{W})$ time, resulting in $\mathcal{O}(\mathcal{W}PN)$ in total for all beads.
Our goal then is to evaluate the probabilities $\Pr\left(\mathbf{w}_{\ell}^{j}\right)$ without explicitly summing over the windings of the other beads. Fortunately, in Section VI of the SI, we show that the probability $\Pr({\{{\mathbf{w}}\}})$ decomposes into
$\Pr({\{{\mathbf{w}}\}})=\prod_{\ell=1}^{N}{\prod_{j=1}^{P}{\Pr\left(\mathbf{w}_{\ell}^{j}\right)}},$ (22)
and
$\Pr\left(\mathbf{w}_{\ell}^{j}\right)=\frac{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1},\mathbf{w}_{\ell}^{j})}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{{j}+1})}.$ (23)
Through Equation 23, it is possible to evaluate each probability in $\mathcal{O}(\mathcal{W})$ time, and all them in $\mathcal{O}(\mathcal{W}PN)$ time in total. As mentioned above, after this step, the forces are evaluated in additional $\mathcal{O}(\mathcal{W}PN)$ time through Equation 21, resulting in $\mathcal{O}(\mathcal{W}PN)$ for the evaluation of the force overall. In more than one dimension, the sum over the windings in Equation 21 splits into separate sums over the coordinates, while Equation 23 splits into a product over the different coordinates; hence it is $\mathcal{O}(\mathcal{W})$ and not $\mathcal{O}(\mathcal{W}^{d})$ (see Section I.D of the SI). Overall the scaling is $\mathcal{O}(\mathcal{W}PN)$ for the force on all beads.
Thus far, the main observation we used to reduce the scaling was the independence of windings of different springs, as apparent in Equation 22 and Figure 2. When we turn to bosons, this is no longer the case. In each permutation separately, the same independence property of windings holds; however, the probability of a permutation depends on all windings, coupling them all. This is the main challenge of the next section.
III.2 Quadratic scaling of bosonic PIMD with PBC
Building on the techniques for including periodic boundary conditions in distinguishable PIMD, we now proceed to describe an efficient algorithm for the bosonic case, achieving quadratic scaling.
To sample the bosonic partition function of Equation 6, similarly to the previous bosonic algorithm (see Section II.2), we use an effective ring-polymer potential $V_{\text{PBC}}^{[{1},{N}]}$ defined by the recurrence relation
$e^{-\beta{V_{\text{PBC}}^{[{1},{N}]}}}$ $=\frac{1}{N}\sum_{k=1}^{N}{e^{-\beta{\left(V_{\text{PBC}}^{[{1},{N-k}]}+E_{\text{PBC}}^{[{{N}-{k}+1},{{N}}]}\right)}}}.$ (24)
The recursion is terminated by setting $V_{\text{PBC}}^{[{1},{0}]}=0$. The crucial difference between Equation 24 and Equation 4, is that the term $E_{\text{PBC}}^{[{{N}-{k}+1},{{N}}]}$ includes a sum over the windings,
$e^{-\beta{E_{\text{PBC}}^{[{u},{v}]}}}$ =$=\sum_{\{{\mathbf{w}}\}^{[{u},{v}]}}e^{-\beta{\frac{1}{2}m\omega_{P}^{2}\sum\limits_{\ell=u}^{v}{\sum\limits_{j=1}^{P}{\left({\bf r}_{\ell}^{j}+\mathbf{w}_{\ell}^{j}L-{\bf r}_{\ell}^{j+1}\right)^{2}}}}}$ (25)
=$=\sum_{\{{\mathbf{w}}\}^{[{u},{v}]}}\prod_{\ell=u}^{v}{\prod_{j=1}^{P}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{j+1},\mathbf{w}_{\ell}^{j})}},$
where $\{{\mathbf{w}}\}^{[{u},{v}]}$ represents the winding vectors of all the beads of particles $u,\ldots,v$. In Equation 25, ${\bf r}_{\ell}^{P+1}={\bf r}_{\ell+1}^{1}$ except ${\bf r}_{v}^{P+1}={\bf r}_{u}^{1}$. Following the existing terminology Feldman and Hirshberg (2023), we call $E_{\text{PBC}}^{[{u},{v}]}$ the cycle energies. By the same reasoning as in distinguishable particles (Equation 15), we can write the cycle energy as
$e^{-\beta{E_{\text{PBC}}^{[{u},{v}]}}}$ $=\prod_{\ell=u}^{v}{\prod_{j=1}^{P}{\mu({\bf r}_{\ell}^{j},{\bf r}_{\ell}^{j+1})}}.$ (26)
We show in Section VI of the SI that the potential of Equation 24 leads to sampling of the correct bosonic partition function with PBC (Equation 6). The remainder of this section explains how we efficiently compute the potential $V_{\text{PBC}}^{[{1},{N}]}$ and the forces it induces, $-\nabla_{{\bf r}_{\ell}^{j}}{V_{\text{PBC}}^{[{1},{N}]}}$ on all the beads. We note that when $\mathcal{W}=0$, Equation 25 coincides with the previous algorithm.
III.2.1 Computing the potential in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ time
We start with evaluating the potential $V_{\text{PBC}}^{[{1},{N}]}$ (Equation 24). The first and most significant step is to evaluate the cycle energies. As in the previous algorithm Feldman and Hirshberg (2023), quadratic scaling is achieved by extending cycles one particle at a time. This is done by adding and removing springs from the ring polymer, but we must include all the possible windings associated with the spring, which is achieved by multiplying and dividing by their statistical weights.
We compute the cycle energies by
$e^{-\beta{E_{\text{PBC}}^{[{u},{v}]}}}=e^{-\beta{E_{\text{PBC}}^{[{u+1},{v}]}}}$ $/\mu({\bf r}_{v}^{P},{\bf r}_{u+1}^{1})$ (27)
$\cdot\mu({\bf r}_{u}^{P},{\bf r}_{u+1}^{1})$
$\cdot e^{-\beta{E_{\textit{int}}^{({u})}}}$
$\cdot\mu({\bf r}_{v}^{P},{\bf r}_{u}^{1})$
In Equation 27, we first remove the contribution of all the windings of the spring that closes the cycle from the last bead of $v$ to the first bead of $u+1$ (see the red dashed lines in the left column of Figure 3, which depicts the three particle case).
Figure 3: Illustration of the idea behind Equation 27, in the case of three bosons and $P=2$, for $u=1$ and $v=3$. For simplicity, we focus only on the exterior spring when presenting the different winding configurations for a cutoff of $\mathcal{W}=1$. The left column depicts some of the winding configurations that contribute to $E_{\text{PBC}}^{[{1},{1}]}$ and $E_{\text{PBC}}^{[{2},{3}]}$. The red and black dashed lines indicate springs that must be removed. The red and cyan solid lines represent the springs that must be added. In the left column, the separate cycles are highlighted in different colors for emphasis.
We then add all the windings of the spring that connects the last bead of the new particle $u$ to the beginning of the previous cycle, the first bead of particle $u+1$ (corresponding to cyan solid lines in the right column of Figure 3). Then, we add the windings of all the interior springs of $u$, as we explain below (not shown in Figure 3). Finally, we add all the windings of the spring that now closes the cycle from the last bead of particle $v$ to the first bead of $u$ (red solid line in the right column of Figure 3).
The contribution of the windings of the interior springs, $e^{-\beta{E_{\textit{int}}^{({u})}}}$, is defined by
$e^{-\beta{E_{\textit{int}}^{({u})}}}=\prod_{j=1}^{P-1}{\mu({\bf r}_{u}^{j},{\bf r}_{u}^{j+1})},$ (28)
which includes all the springs of particle $u$ that do not depend on the connectivity of $u$ in the cycle. The recurrence of Equation 27 is terminated by cycles of just one particle:
$e^{-\beta{E_{\text{PBC}}^{[{u},{u}]}}}=\mu({\bf r}_{u}^{1},{\bf r}_{u}^{P})\cdot e^{-\beta{E_{\textit{int}}^{({u})}}}.$ (29)
The manipulations performed in Equation 27 rely on the fact that it is possible to sum over the winding vectors of springs of each cycle separately.
With these expressions, evaluating all the cycle energies requires $\mathcal{O}(\mathcal{W}(N^{2}+PN))$. The most time consuming part is evaluating Equation 16 for $\mathcal{O}(PN)$ internal springs and $\mathcal{O}(N^{2})$ springs between the last and the first bead of each pair of particles, and each such evaluation requires $\mathcal{O}(\mathcal{W})$ time. As mentioned above, even in more than one dimension, the scaling is $\mathcal{O}(\mathcal{W})$ and not $\mathcal{O}(\mathcal{W}^{d})$ (see Section I.D of the SI). Once the statistical weights are known, evaluating the contribution of the interior springs according to Equation 28 takes $\mathcal{O}(P)$ time for each of the $N$ particles; evaluating Equation 29 is $\mathcal{O}(1)$ for each of the $N$ single-particle cycles; using Equation 27 takes additional $\mathcal{O}(1)$ for each of the $\mathcal{O}(N^{2})$ cycles. After evaluating the cycle energies, evaluating the potentials $V_{\text{PBC}}^{[{1},{1}]},\ldots,V_{\text{PBC}}^{[{1},{N}]}$ of Equation 24 takes only additional $\mathcal{O}(N^{2})$ time ($\mathcal{O}(N)$ for each potential). Thus the potential $V_{\text{PBC}}^{[{1},{N}]}$ is computed in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ overall. We now turn to computing the forces, in quadratic time as well.
III.2.2 Computing the force in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ time
To compute the force induced by $V_{\text{PBC}}^{[{1},{N}]}$, we first express the force as a weighted average over the configurations stemming from different permutations and windings. In Section VI of the SI, we show that $V_{\text{PBC}}^{[{1},{N}]}$ can be written as
$e^{-\beta{V_{\text{PBC}}^{[{1},{N}]}}}=\frac{1}{{N}!}\sum_{\sigma}{\sum_{\{{\mathbf{w}}\}}{e^{-\beta{E^{G[{\sigma}],\{{\mathbf{w}}\}}}}}},$ (30)
and that this is equivalent to Equation 24.
This is similar to the previous bosonic algorithm, except that the choice of winding vectors generates more configurations for every permutation. In Equation 30, the function $G$ replaces some permutations by others, because, as in the previous algorithm, not all permutations directly appear in $V_{\text{PBC}}^{[{1},{N}]}$; the definition is the same as in the previous algorithm, and appears in Section I.C of the SI.
By taking the derivative of Equation 30, the force can be written as a weighted average over permutations and windings,
$-\nabla_{{\bf r}_{\ell}^{j}}{V_{\text{PBC}}^{[{1},{N}]}}=\sum_{\sigma}{\sum_{\{{\mathbf{w}}\}}{\Pr(G[{\sigma}],{\{{\mathbf{w}}\}})\cdot-\nabla_{{\bf r}_{\ell}^{j}}{E^{G[{\sigma}],\{{\mathbf{w}}\}}}}},$ (31)
where
$\Pr(G[{\sigma}],{\{{\mathbf{w}}\}})=\frac{e^{-\beta{E^{G[{\sigma}],\{{\mathbf{w}}\}}}}}{{N}!\cdot e^{-\beta{V_{\text{PBC}}^{[{1},{N}]}}}},$ (32)
is a Boltzmann distribution over the configurations. The force in a configuration $-\nabla_{{\bf r}_{\ell}^{j}}{E^{G[{\sigma}],\{{\mathbf{w}}\}}}$ is the same as in Equation 19, except using ${\bf r}_{\ell}^{P+1}={\bf r}_{\sigma(\ell)}^{1}$, according to the permutation.
Computing the force on a bead efficiently is based on grouping the configurations that exert the same force on the bead. As in the previous bosonic algorithm, we separate the force evaluation for the first and last beads of each particle, referred to as exterior beads, and the rest of the beads, which are referred to as interior beads.
Force on interior beads
The force on bead $j\neq 1,P$ of particle $\ell$ is the same as in the distinguishable particle case with PBC (Equation 21). The reason is that the force on an interior bead ${\bf r}_{\ell}^{j}$ in each configuration is independent of the permutation $\sigma$ and only depends on the winding vectors $\mathbf{w}_{\ell}^{j},\mathbf{w}_{\ell}^{j-1}$. Thus the expression reduces to a sum over winding vectors, which can be treated the same way as in distinguishable particles. These forces are evaluated in $\mathcal{O}(\mathcal{W}PN)$ for all the $P-2$ interior beads of each of the $N$ particles.
Force on exterior beads
The force on beads $1,P$ of each particle depends both on the permutation and on winding (see Figure 1). Our approach is similar to the distinguishable case (Equation 21), grouping contributions that exert the same force on a given bead. Accounting for both permutations and windings leads to the following expression for bead $P$,
$-\nabla_{{\bf r}_{\ell}^{P}}{V_{\text{PBC}}^{[{1},{N}]}}=$ (33)
$\sum_{\mathbf{w}_{\ell}^{P-1}}{\Pr\left(\mathbf{w}_{\ell}^{P-1}\right)\cdot-m\omega_{P}^{2}({\bf r}_{\ell}^{P}-{\bf r}_{\ell}^{P-1}-\mathbf{w}_{\ell}^{P-1}L)}$
$+$ $\sum_{{\ell}^{\prime}=1}^{N}{\sum_{\mathbf{w}_{\ell}^{P}}{\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}},\mathbf{w}_{\ell}^{P}\right)\cdot-m\omega_{P}^{2}({\bf r}_{\ell}^{P}+\mathbf{w}_{\ell}^{P}L-{\bf r}_{{\ell}^{\prime}}^{1})}}.$
Equation 33 contains contributions from two springs: the spring connecting bead $P$ to the previous bead, ${\bf r}_{\ell}^{P-1}$, and the spring connecting it to the next bead, ${\bf r}_{G[{\sigma}](\ell)}^{1}$, which depends on the permutation. The first term averages over all the winding vectors for the former, and the second over all the winding vectors of the latter, as well as permutations. An analogous expression for the force on the first bead appears in Section I.A of the SI.
The probability $\Pr\left(\mathbf{w}_{\ell}^{P-1}\right)$ for the spring that is not affected by exchange is the same as in Equation 23. The other spring, however, is affected by exchange, and so Equation 23 does not apply. Instead, Equation 33 uses the joint probability $\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}},\mathbf{w}_{\ell}^{P}\right)$. It is defined as the sum of the probabilities of all the configurations where the permutation satisfies $G[{\sigma}](\ell)={\ell}^{\prime}$ and the winding vector of bead $P$ of particle $\ell$ is equal to $\mathbf{w}_{\ell}^{P}$. This joint probability can be written as a product,
$\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}},\mathbf{w}_{\ell}^{P}\right)=$ (34)
$\quad\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)\cdot\Pr\left(\mathbf{w}_{\ell}^{P}\mid{G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right),$
where $\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)$ is the probability that bead $P$ of particle $\ell$ is connected to the first bead of particle ${\ell}^{\prime}$, and $\Pr\left(\mathbf{w}_{\ell}^{P}\mid{G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)$ is the conditional probability of having $\mathbf{w}_{\ell}^{P}$ given that $G[{\sigma}](\ell)={\ell}^{\prime}$. The benefit of this decomposition is that both these probabilities can be evaluated efficiently, as we now show.
Connection probability
The marginal probabilities $\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)$ retain the same form as in the previous algorithm Feldman and Hirshberg (2023), the only difference being that the term $E_{\text{PBC}}^{[{u},{v}]}$ includes the sum over windings per Equation 25. For instance, we show in Section VI of the SI that for $\ell^{\prime}\leq\ell$,
$\Pr\left({G[{\sigma}]}({\ell})={\ell^{\prime}}\right)=\frac{1}{\ell}\frac{e^{-\beta{\left(V_{\text{PBC}}^{[{1},{\ell^{\prime}-1}]}+E_{\text{PBC}}^{[{\ell^{\prime}},{\ell}]}+V_{\text{PBC}}^{[{\ell+1},{N}]}\right)}}}{e^{-\beta{V_{\text{PBC}}^{[{1},{N}]}}}},$ (35)
where the partial bosonic potentials are defined through the recurrence relation
$e^{-\beta{V_{\text{PBC}}^{[{u},{N}]}}}=\sum_{\ell=u}^{N}{\frac{1}{\ell}e^{-\beta{\left(E_{\text{PBC}}^{[{u},{\ell}]}+V_{\text{PBC}}^{[{\ell+1},{N}]}\right)}}},$ (36)
with $u=1,\dots,N$. Expressions for all the connection probabilities appear in Section I.B of the SI.
The derivation of Equations 35 and 36 relies of the fact that a permutation can be decomposed into disjoint cycles. With periodic boundary conditions, the sum over permutations and windings decomposes too into a sum over disjoint cycles and windings. This is the reason that the connection probabilities have the same expression as in the previous algorithm. In fact, this is what allows us to define the potential with PBC itself in a recursive manner (Equation 24).
Conditional winding probability
The conditional probability is defined as a sum over all the permutations in which $G[{\sigma}](\ell)=\ell^{\prime}$ and have the winding $\mathbf{w}_{\ell}^{P}$. We show in Section VI of the SI that it can be calculated efficiently by
$\Pr\left(\mathbf{w}_{\ell}^{P}\mid{G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)=\frac{\mu({\bf r}_{\ell}^{P},{\bf r}_{{\ell}^{\prime}}^{1},\mathbf{w}_{\ell}^{P})}{\mu({\bf r}_{\ell}^{P},{\bf r}_{{\ell}^{\prime}}^{1})}.$ (37)
This expression is identical to Equation 23, except that here ${\bf r}_{\ell}^{P+1}={\bf r}_{{\ell}^{\prime}}^{1}$. As Equation 37 shows, once the connectivity is set, all permutations have the same probability of winding $\mathbf{w}_{\ell}^{P}$, which has the same form as in the case of distinguishable particles.
Complexity of force evaluation
Overall, the force on all the exterior beads is computed in $\mathcal{O}(\mathcal{W}N^{2})$ time: first, the connection probabilities $\Pr\left({G[{\sigma}]}({\ell})={{\ell}^{\prime}}\right)$ are computed in $\mathcal{O}(N^{2})$ according to Equation 35 and the rest of the expressions appearing in Section I.B of the SI. Second, the $\mathcal{W}N^{2}$ joint probabilities are computed in $\mathcal{O}(1)$ each according to Equation 37. Finally, the force on each of the $2N$ exterior beads is computed in $\mathcal{O}(\mathcal{W}N)$ according to Equation 33. The same considerations apply when evaluating the force on the first bead of each particle. As before, in the multidimensional case, the sum over windings in Equation 33 splits into separate sums for different axes, hence it is $\mathcal{O}(\mathcal{W})$ and not $\mathcal{O}(\mathcal{W}^{d})$ (see Section I.D of the SI).
III.2.3 Estimator for the kinetic energy in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ time
We derive a thermodynamic kinetic energy estimator appropriate for PBC, which is given by
$\expectationvalue{E}=\frac{dPN}{2\beta}+\expectationvalue{V_{\text{PBC}}^{[{1},{N}]}+\beta\frac{\partial V_{\text{PBC}}^{[{1},{N}]}}{\partial\beta}}+\expectationvalue{\bar{U}},$ (38)
where the brackets denote an average over the ensemble with PBC. This is the same estimator as in the previous algorithm, except that the recurrence relation for $V_{\text{PBC}}^{[{1},{N}]}+\beta\frac{\partial V_{\text{PBC}}^{[{1},{N}]}}{\partial\beta}$ must include the cycle energies with the sum over windings, $E_{\text{PBC}}^{[{u},{v}]}$, and their derivative with respect to the inverse temperature $\beta$. We provide full details on how this expression is evaluated efficiently in Section II of the SI.
III.2.4 Additional implementation details
Pseudocode for our algorithm appears in Section III of the SI. In our implementation, we also take care of the following:
Wrapping coordinates
The limits of integration in Equation 6 restrict the coordinates of the particles to the unit cell. Therefore, in the course of the simulation, we prevent the particles from leaving the box by adding or subtracting integer multiples of the box length, whenever the particle leaves, as is customary in MD simulations with PBC Allen and Tildesley (2017).
Numerical stability
The log-sum-exp method Blanchard et al. (2020) is used Xiong and Xiong (2022a) to ensure numerical stability whenever a sum of exponentials is used. In addition to the stages where such a sum is performed in the previous algorithm, our algorithm also includes such a sum as part of computing the cycle energies (described in Equation 27).
III.3 Numerical results
We applied our algorithm in PIMD simulations of two periodic systems: the free Bose gas, and particles in a sinusoidal trap. Below, we benchmark the new algorithm (labeled as “pimd-b (pbc)”) and compare it to analytical results (labeled as “exact (pbc)”), and simulations which neglect PBC using the previous algorithm (labeled as “pimd-b (no pbc)”). Then, we validate the theoretical scaling of the algorithm, quadratic with $N$ and linear with $\mathcal{W}$.
Benchmark
In the free Bose gas, Figure 4 presents the resulting energy from our bosonic PIMD with PBC algorithm, at a range of temperatures, for $N=64$ and a density of $0.035\;\text{\AA}^{-3}$. A maximum of $P=32$ beads and a winding cutoff of $\mathcal{W}=1$ were required for convergence. Additional details appear in Section IV of the SI. For comparison, the results of the previous algorithm are also presented. The analytical solution for the free Bose gas with PBC is derived in Section V of the SI. We find very good agreement between our results and the analytical solution, with a relative error between 0.5% and 3.3%, in a temperature range where neglecting PBC leads to large errors.
Figure 4: Energy per particle as a function of the temperature for the free Bose gas with $N=64$ atoms and density $n=0.035\;\text{\AA}^{-3}$. The data points correspond to bosonic PIMD simulations with PBC (blue) and without PBC (green). The solid line is the analytical result with PBC. The bottom panel shows absolute energy values, while the top panel shows the relative error of the PIMD-B PBC results when compared to the exact results.
Figure 5: Energy per particle as a function of the temperature of $N=32$ particles in a sinusoidal trap. The data points correspond to bosonic PIMD simulations with PBC (blue) and without PBC (green). The solid line is the analytical result for the periodic system. The bottom panel shows absolute values, while the top panel shows the relative error of the PBC results when compared to the exact result.
For the periodic sinusoidal trap, Figure 5 presents the energy of a system of $N=32$ particles at the same density. Here, too, our algorithm reproduces the analytical result excellently, with a relative error ranging from $0.03\%$ to $1.3\%$. See Section V of the SI for the derivation of the analytical result. For comparison, PIMD simulations with the previous algorithm are also presented, and lead to large errors. These results demonstrate the correctness of our method of including PBC in bosonic PIMD.
Scaling
Figure 6 shows the time required for a PIMD step with our bosonic algorithm as a function of $N$ in a simulation of free particles and $\mathcal{W}=1$ (blue). The results are consistent with quadratic scaling with $N$, with a slope of $1.85$ in a log-log scale and a Pearson correlation coefficient $R^{2}=0.9998$. Including PBC with $\mathcal{W}=1$ is roughly ×$\times 10$ slower in comparison to the previous algorithm (green).
Figure 7 shows the time required for a PIMD step with our algorithm in a simulation of $N=64$ free particles and a varying winding cutoff $\mathcal{W}$. The results are consistent with linear scaling with $\mathcal{W}$ in a log-log scale and a Pearson correlation coefficient $R^{2}=0.9994$. In Figures 6 and 7, the time measured in each point is averaged over $1000$ steps, the temperature is $T=3.0\;\mathrm{K}$, and $P=32$ and $P=4$ for for Figure 6 and Figure 7, respectively.
Figure 6: Scaling with the number of particles $N$ of bosonic PIMD with PBC and $\mathcal{W}=1$ (blue). For comparison, we also show the scaling of bosonic PIMD with the minimum-image convention, as explained in Section IV (orange). Both algorithms are compared to the original non-periodic bosonic algorithm (green). In all cases,the fitted slope in log-log scale is close to $2$, as expected for quadratic scaling, but the prefactor varies.
Figure 7: Scaling of bosonic PIMD with PBC as a function of the winding cutoff $\mathcal{W}$. In log-log scale, the fitted slope is close to $1$, as expected for linear scaling.
IV PBC vs. the minimum image convention
In this section, we compare the new algorithm for bosonic PIMD with PBC to applying the minimum-image convention (MIC) to the springs as an approximate way of including PBC in the simulation. Specifically, we use the previous algorithm without PBC, but 1) wrap coordinates inside the unit cell, and 2) replace the differences ${\bf r}_{\ell}^{j}-{\bf r}_{\ell}^{j+1}$ in the spring energies and forces with the displacement to the nearest periodic image, by choosing the winding vector $(\mathbf{w}_{\ell}^{j})^{\ast}$ that minimizes it,
$(\mathbf{w}_{\ell}^{j})^{\ast}=\mathop{\text{arg min}}_{\mathbf{w}_{\ell}^{j}\in\mathbb{Z}^{d}}\{{{\bf r}_{\ell}^{j}-{\bf r}_{\ell}^{j+1}+\mathbf{w}_{\ell}^{j}L}\}.$ (39)
In the multidimensional case each scalar component of Equation 39 is minimized separately. In essence, the MIC selects a single winding vector per bead for which the energy is minimal, and the probability is the largest according to Equations 37 and 23. Although very reasonable, this approach is not rigorous, since the MIC algorithm does not take into account the exponentially-many other choices of winding vectors. We next investigate when this strategy works or fails in practice.
We find that in the periodic systems we consider, the MIC approximation converges to the correct result when $P$ is sufficiently large. However, we find that this $P$ can be unnecessarily large compared to the convergence of the rigorous bosonic PIMD with PBC algorithm in some cases. We demonstrate that the difference between the algorithm with PBC and the MIC approximation disappears for $P$ that is large enough so that the distribution of winding numbers is narrowly centered around a single value. This explains why the MIC approximations converges to the correct result in practice, and why it fails to do so in values of $P$ lower than required for convergence of the PBC algorithm.
MIC approximation validity
Figure 8 shows the convergence with $P$ of both the algorithm with PBC and the MIC approximation for the free Bose gas with $N=64$, $T=0.5\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. If not shown, the statistical error is smaller than the symbol size. We find that MIC converges to the correct result with PBC at about $P=20$. Figure 9 shows the convergence with $P$ for bosons in a sinusoidal trap with $N=32$, $T=1.0\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. We find that the MIC approximation converges to the correct result with PBC at about $P=32$ with the algorithm the rigorously accounts for PBC is converged already for $P=4$. As another example,  Figure 9 compares the results of rigorous PBC treatment with the MIC for the sinusoidal trap. There is a small difference between the PBC algorithm and the MIC approximation in this case, and both converge at about $P=26$.
Criterion for using the MIC approximation
Figure 8: Energy per particle as a function of the number of beads $P$ for $N=64$ free bosons at $T=0.5\;\mathrm{K}$ and a density of $n=0.035\;\text{\AA}^{-3}$. The PBC algorithm converges at $P=4$ while the MIC approximation requires about $P=14$ beads.
Figure 9: Energy per particle as a function of the number of beads $P$ for $N=32$ bosons in a sinusoidal trap at $T=1.0\;\mathrm{K}$ and density $n=0.035\;\text{\AA}^{-3}$. The difference between the PBC algorithm and MIC is negligible and both converge at the same $P$.
To understand when MIC is a good approximation , we consider for each spring the quantity
$1-\Pr\left((\mathbf{w}_{\ell}^{j})^{\ast}\right),$ (40)
where $(\mathbf{w}_{\ell}^{j})^{\ast}$ is the winding vector chosen by MIC accroding to Equation 39 for bead $j$ of particle $\ell$. Equation 40 is the total probability of all the winding vectors neglected by the MIC. Notice that for $j=P$, $\Pr\left((\mathbf{w}_{\ell}^{P})^{\ast}\right)$ is a weighted sum over the possible connections of this bead. We now show that this quantity is indicative of whether MIC is a good approximation to the rigorous PBC treatment.
Figure 10 shows the discarded probability for bead $j=P$ of particle $\ell=1$ in a simulation of the free Bose gas at the same conditions as Figure 8. When $P=4$, the discarded probability is high during the simulation, i.e., there are multiple winding vectors with comparable, non-negligible probabilities, and the choice of a single winding vector by MIC neglects a significant share of important winding configurations. As $P$ increases, the discarded probability is lower throughout the simulation, and for $P=14$ it is nearly negligible throughout the entire simulation. These results match Figure 8, where the MIC converges at $P=14$, and differs significantly from the PBC results, which were already converged at $P=4$.
Figure 11 shows the discarded probability for the same bead but for the sinusoidal trap at the conditions of Figure 9. The differences between MIC and PBC are relatively small, which is also evident from the discarded configurations. In $P=4$, the discarded probability is non-negligible but small, and indeed there are small differences for $P=4$ in Figure 8. When the discarded probability becomes negligible, as $P$ increases to $8$ and $14$, the small differences in energy disappear, in agreement with Figure 8.
The reason that the discarded probability decreases when $P$ increases can be understood from the probability distribution of winding numbers per Equations 23 and 34. Larger $P$ makes the springs stiffer, leading to a higher energetic penalty for stretching them. This leads to a narrower winding distribution, and to MIC converging similarly to the rigorous treatment of PBC. The same considerations apply when the box length increases, as a larger $L$ more strongly penalizes stretched springs.
Figure 10: The probability of all the winding configurations for ${\bf r}_{1}^{P}$ neglected by the MIC approximation during a simulation of $N=64$ free bosons at $T=0.5\;\mathrm{K}$, for different values of $P$. As $P$ increases, the probability decays and the MIC approximation becomes better. At $P=14$ MIC and Winding yield practically identical trajectories, which is also when $\expectationvalue{E}/N$ in a MIC simulation converges to the correct result (see Figure 8).
Figure 11: The probability of all the winding configurations neglected by the MIC approximation during a simulation of $N=32$ bosons in a sinusoidal trap at $T=1.0\;\mathrm{K}$, for values of $P$. As $P$ increases, the probability decays and the MIC approximation becomes better. At $P=14$ MIC and Winding yield practically identical trajectories, which is also when $\expectationvalue{E}/N$ in a MIC simulation converges to the correct result (see Figure 9).
V Summary and conclusions
In this paper, we developed an algorithm that rigorously accounts for PBC in bosonic PIMD simulations. It required summing over the spring interaction between neighboring beads over all periodic images. The key difficulty stemmed from the exponential number of periodic images (winding vectors) that had to be included, and that the winding vector of the last bead of every particle was coupled with the permutations, of which there is also an exponential number. We recently developed a quadratic scaling, recursive algorithm for handling the permutations but did not account for the sum over windings. Here, we showed that we can maintain the same recursive structure of the bosonic ring polymer potential and rigorously account for PBC by including the sum over windings in the cycle energies. We also evaluated the forces efficiently by deriving expressions for the connection and winding probabilities. The resulting algorithm scales quadratically with the number of particles and linearly with the winding cutoff $\mathcal{W}$. We used the new algorithm to perform bosonic PIMD simulations with rigorous treatment of PBC for the first time, and benchmarked our approach on the free Bose gas and a model of cold atoms in a sinusoidal trap. Finally, we carefully examined an approximate treatment of PBC using the MIC, derived a quantitative criterion to when is is a good approximation to the full periodic algorithm, and provided examples. We found that the MIC approximation becomes increasingly more suitable when the number of beads increases. Our works enables simulations of bosonic systems with PIMD while accurately accounting for PBC.
Acknowledgements.
B.H. acknowledges support by the USA-Israel Binational Science Foundation (grant No. 2020083) and the Israel Science Foundation (grants No. 1037/22 and 1312/22). Y.F. was supported by Schmidt Science Fellows, in partnership with the Rhodes Trust.